OpenStudy (anonymous):

Solve |x + 6| = 12. A. 6 B. -18 C. {6, -18} D. No Solution

OpenStudy (turingtest):

we get two equations when we have an absolute value, because there are two possibilites

OpenStudy (turingtest):

\[|x+6|=12\]implies that either\[x+6=12\]or\[x+6=-12\]

OpenStudy (anonymous):

-12=x+6 x+6=12 -12-6=x x=12-6 -18=x x=8

OpenStudy (turingtest):

solve each equation separately and you will have your answer

OpenStudy (anonymous):

@TuringTest I'm having a doubt ... example: if I have this equation: \[\Large |x-2|>7\] when I start solving, after taking off the absolute value should I change - into + like: -7>x+2>7 ?? :(:( or do I have to leave it minus like it is ?

OpenStudy (turingtest):

@Kreshnik 1)please don't cut me off in the future, I had this under control and wanted to \(not\) provide the answer -in answer to your question \[|x-2|>7\]implies that either \[x-2>7\]or\[x-2<-7\]

OpenStudy (turingtest):

only if the sign is\[x<y\]can you write\[-y<x<y\](provided that y>0}

OpenStudy (turingtest):

sorry, typo* should have been\[|x|\le y\implies -y\le x\le y\]

OpenStudy (anonymous):

... and if \[\Large x\neq y\] x=1 y=2 1<2 -2<1<2 I think yes :( @TuringTest I do really apologize for interrupting , It won't happen again !

OpenStudy (turingtest):

but\[|x|\ge y\implies x\ge y\text{ or } x\le-y\]

OpenStudy (turingtest):

it's fine, no worries ;D I'm just trying to prevent those who would keep trying to avoid explanation, and I think this particular asker is a culprit ;)

OpenStudy (anonymous):

ahh... obviously, he doesn't even try to solve those, and I think that will revenge him in the future! ... he won't get nowhere without trying ! ... @TuringTest Thanks for explaining it. I apologize again ;)

OpenStudy (anonymous):

Yup, I think so @TuringTest and @kreshnik - I'll take care of that. Btw, great work from both of you here.